OpenStudy (anonymous):

The area of a rectangular rug in Monica’s living room is 202.5 square feet. If the length of the rug is 16.2 feet, what is its width?

OpenStudy (bibby):

The area of a rectangle is l*w. So we have: l=16.2 A=l*w=202.2 Solve for w

OpenStudy (calculusxy):

\[16.2 \times w = 202.5\]

OpenStudy (calculusxy):

To solve for w, you need to do the inverse of multiplication, meaning you should divide.

OpenStudy (skullpatrol):